Marigold Corp. had 192000 shares of common stock, 19900 shares of convertible preferred stock, and $1600000 of 5% convertible bonds outstanding during 2018. The preferred stock is convertible into 40100 shares of common stock. During 2015, Marigold paid dividends of $0.90 per share on the common stock and $4 per share on the preferred stock. Each $1,000 bond is convertible into 30 shares of common stock. The net income for 2018 was $597000 and the income tax rate was 40%.

Diluted earnings per share for 2018 is (rounded to the nearest penny)

$2.30.
$2.57.
$2.10.
$2.14.
Answer #1

Bond convert into shares = $1600000*30/1000 = 48000 Shares

Interest expense after tax = 1600000*5%*60% = $48,000

Diluted earning per share = Adjusted net income/Adjusted shares

= ($597000 + 48000)/(192000+40100+48000)

Diluted earning per share = 2.30 per share

Option 1. is correct answer.